### v2

# Illuminating Blind Spots: Exploring LLMs as a Source of Targeted Synthetic Textual Data to Minimize High Confidence Misclassifications

Abstract
Natural Language Processing (NLP) models optimized for predictive performance often make high confidence errors and suffer from vulnerability to adversarial and out-of-distribution data. Existing work has mainly focused on mitigation of such errors using either humans or an automated approach. In this study, we explore the usage of large language models (LLMs) for data augmentation as a potential solution to the issue of NLP models making wrong predictions with high confidence during classification tasks. We compare the effectiveness of synthetic data generated by LLMs with that of human data obtained via the same procedure. For mitigation, humans or LLMs provide natural language characterizations of high confidence misclassifications to generate synthetic data, which are then used to extend the training set. We conduct an extensive evaluation of our approach on three classification tasks and demonstrate its effectiveness in reducing the number of high confidence misclassifications present in the model, all while maintaining the same level of accuracy. Moreover, we find that the cost gap between humans and LLMs surpasses an order of magnitude, as LLMs attain human-like performance while being more scalable. [^1] Code is made available at: see supplementary materials.
Illuminating Blind Spots: Exploring LLMs as a Source of Targeted Synthetic Textual Data to Minimize High Confidence Misclassifications
Philip Lippmann, Matthijs T.J. Spaan, Jie Yang Delft University of Technology p.lippmann@tudelft.nl
1 Introduction
Natural Language Processing (NLP) models with the sole goal of optimizing predictive performance have achieved remarkable accuracy across a wide range of tasks but are often vulnerable to data that is adversarial in nature Papernot et al. (2016); Wang et al. (2019); Brown et al. (2020). Robustness – a model’s ability to deal with diverse and challenging data – plays a crucial role in addressing the vulnerability of NLP models to adversarial and out-of-distribution data Wang et al. (2022); Du et al. (2023), though it may be at odds with accuracy Tsipras et al. (2019).
[…]
Sentiment Prediction: Negative
Confidence: 0.98
[…]
Sentiment Prediction: Positive
Confidence: 0.94   Outcome: UU
[…]
Table 1: Rows 1 & 2 show an attack on a sentiment classifier using a perturbation method. Substitution of the perturbed text leads to change for a correct prediction to a wrong prediction at high confidence, leading to a UU. Row 3 is a generalized hypothesis that describes the cause of the UU. Row 4 is the exploration hypothesis based on the original hypothesis in Row 3.
Figure 1: Procedure performed by human or LLM, beginning from a single UU present in a blind spot (denoted by a cross). This UU is then used to generate an initial hypothesis via abstraction. This abstraction hypothesis can then either by used to generate a synthetic sample that targets the existing blind spot or to generate a new hypothesis via exploration, which in turn is then used to generate a synthetic sample that targets an unknown blind spot.
Prediction errors can be categorized according to the model’s confidence in its prediction, with unknown unknowns (UUs) corresponding to misclassifications where the model is very confident Attenberg et al. (2015). UUs have been shown to cluster to form blind spots, which represent areas in the feature space that lead a model to produce high confidence misclassifications Lakkaraju et al. (2017); Liu et al. (2020). NLP models are prone to produce UUs as a consequence of text perturbations, such as the substitutions of relevant synonyms Ren et al. (2019). In the first two rows of table 1, we show an example of such a perturbation that successfully changes the predicted label, resulting in a UU. Discovering and mitigating blind spots is critical in high-stakes NLP tasks, where their presence can result in unwanted adverse outcomes, such as unreliable suicide prediction models Large et al. (2017) and biased sentencing decisions in criminal justice Crawford (2016).
The identification of UUs has been extensively explored in existing literature Attenberg et al. (2015); Bansal and Weld (2018); Vandenhof (2019); Liu et al. (2020); Han et al. (2021), with more recent researcher involving humans to aid blind spot identification efforts Cabrera et al. (2021); Han et al. (2021). However, current methods incorporate humans in a heuristic manner, not designing the task effectively to fully leverage their generalization abilities for blind spot mitigation. The non-triviality of mitigating blind spots in NLP models lies in the challenge of generalizing from previously identified individual UUs to the mitigation of unseen blind spots. It remains an open question how to best engage humans – or alternative agents, such as large language models (LLMs) – to effectively mitigate blind spots at scale.
To this end, we explore the characterization of existing blind spots by either humans or LLMs as a potential solution to the issue of NLP models making wrong predictions with high confidence, emphasizing the importance of generalization. Generalization refers to the idea that intelligent agents employ prior knowledge for new learning situations that share similarities with past experiences Gluck et al. (2011); Banich and Caccamise (2010); Allaway and McKeown (2020). We pose that by utilizing generalized patterns and similarities between seen and unseen high confidence misclassifications, blind spots can be effectively described by humans or LLMs. To accomplish this, we guide both humans and LLMs in formulating natural language hypotheses that describe the underlying principles contributing to the formation of blind spots. After generating hypotheses based on discovered UUs, we generate relevant additional synthetic samples which target UUs. In our study, we specifically compare the effectiveness of the synthetic data contributed by humans and LLMs, respectively.
We find that our approach is capable of greatly reducing the number of high confidence misclassifcations without decreasing accuracy. Further, we find LLMs exhibit a greater capacity for characterizing blind spots compared to humans, as evidenced by a larger average decrease in the number of UUs via our LLM-based approach (19.54%) compared to our human-based one (16.80%). Finally, the difference in cost per human-generated datum versus LLM-generated datum is substantial, with LLM-generated data being significantly more cost-effective.
2 Related Work
Before outlining our methodology, we introduce relevant prior research for approaches to high confidence misclassifications, as well as how others have tried to avoid such model behaviour.
2.1 Unknown Unknowns
Attenberg et al. (2015) introduce the concept of querying humans to find UUs in a game-like setting and show that there were patterns to the found UUs. Vandenhof (2019) proposes an approach to identify UUs where human-interpretable decision rules are learned to approximate how a model makes high-confidence predictions. Crowdworkers then contradict these rules by finding an instance that would classify as a UU. Han et al. (2021) propose an approach where crowdworkers continuously extend a dataset with additional UUs, that the chosen model is iteratively trained on. Cabrera et al. (2021) explore the use of crowdworkers to generate failure reports for NLP models to describe how or why the model failed. There are also algorithmic approaches to finding UUs, such as Lakkaraju et al. (2017), who propose utilizing an explore-exploit approach to find groups of UUs. Bansal and Weld (2018) extend this by proposing a utility model that rewards the degree to which the found UUs cover a sample distribution, thus encouraging the discovery of new blind spots. Instead, we do not find the UUs algorithmically, but instead use an LLM or crowdworkers to find existing UUs, extrapolate from these to unseen UUs, and generate synthetic data targeting these.
2.2 Model Calibration and Robust Training
The concept of UUs and blind spots is connected to model calibration (Guo et al., 2017; Minderer et al., 2021; Tian et al., 2023). A model that is well-calibrated will have its prediction confidence aligned with the likelihood of the correctness of the prediction. Investigating a model with blind spots is an examination of a model that is not well-calibrated. Therefore, our approach improves the calibration of the model.
In the case where the UUs are specifically generated through adversarial attacks, illumination of model blind spots is also related to robust training. UUs that populate these blind spots, when created by such attacks, may be identified as adversarial examples (Ribeiro et al., 2018; Wallace et al., 2019; Wang et al., 2020). This underscores the relationship between our proposed method and robust training practices, as well as the creation and identification of adversarial examples through adversarial attacks, with the aim of improving the robustness of the model (Madry et al., 2018; Pang et al., 2021). Unlike general robust training practices, our method focuses not on general robustness – but rather on high confidence misclassifications – and is not limited to just adversarial samples, as we also consider UUs that occur naturally, without any perturbation.
For the NLP research area, several approaches have been proposed to utilize synthetic data to expand training sets (Puri et al., 2020; Claveau et al., 2021). Synthetic textual data has previously been proposed specifically for sentiment classification tasks (Maqsud, 2015).  He et al. (2022) explore few-shot prompting LLMs to generate task specific synthetic training data. Unlike prior work, we propose a method to generate targeted synthetic data with the purpose of eliminating blind spots that lead to high confidence misclassifications.
3 Methodology
In our proposed approach, for the purpose of blind spot mitigation, a human or LLM is asked to perform one of three possible tasks: hypothesis generation via abstraction, hypothesis generation via exploration, or sample generation. These are described below and shown schematically in figure 1.
3.1 Problem Formulation
For passive UU discovery, let the dataset be $\mathcal{D}=\{(x_{1},y_{1}),...,(x_{n},y_{n})\}$, where $x$ is the original text sample and $y$ the original label. Without having access to $y$, a predictive model $\theta$ is used to obtain $y=\theta(x)$ at a confidence $c\in[0,1]$. Formally, a UU occurs when (1) the wrong label is predicted by $\theta$ and (2) the prediction is made at $c\geq\tau$, indicating high confidence. In the case of proactive UU discovery, as is performed in this work, an additional technique is used to increase the misclassifications produced by $\theta$. For this, a black-box adversarial perturbation model $G$ generates perturbed samples $x^{\prime}=G(x)$, where $x^{\prime}\neq x$. Then $\theta$ is used to predict the new label for every sample $y^{\prime}=\theta(x^{\prime})$ at $c\in[0,1]$. The newly formed perturbed dataset $\mathcal{P}$ constitutes of these new labels and their corresponding $x^{\prime}$. A perturbation may be considered successful if $y^{\prime}\neq y$. If a perturbation occurs, there is an additional requirement for a UU: (3) $x^{\prime}$, regardless of its label indicated by the classifier, maintains the same underlying true label as $x$ post perturbation.
3.2 Generalization via Hypothesis Creation
For generalization, we present humans or LLMs with related pairs from $(x,y)\sim\mathcal{D}$ and $(x^{\prime},y^{\prime})\sim\mathcal{P}$ consisting of UUs and ask them to create hypotheses regarding the underlying causes of present blind spots. Due to the ability of humans to learn and generalize effectively using sparse amounts of data Lake et al. (2015), we can focus on subsets of UUs as relevant samples for the hypothesis generation task. In our study, hypotheses take the form of simple sentences that describe why a UU occurred and what information can be generalized from it to other potential UUs. As such, it does not act as a simple explanation of the failure case. Any hypothesis should cover multiple sentences indiscriminately, should the same characteristics persist across them. This allows a hypothesis to address multiple UUs that cluster together to form a blind spot and thus illuminate it. Examples of what these hypotheses might look like in practice are shown in the final two rows of table 1.
Abstraction Abstraction constitutes the usage of a single hypothesis to describe part of, or ideally, an entire blind spot, to expand the knowledge regarding that particular blind spot. During the abstraction step, we present UUs to both human participants and the LLM. They are shown a sample $(x_{i},y_{i})$ and its perturbed form $(x_{i}^{\prime},y_{i}^{\prime})$, and are tasked with generating a hypothesis that explains any factors that could have led to it being a UU. We frame the task in a way that this could be related to a pattern, semantics, syntax, specific words, or something else between the original and perturbed samples. They are tasked with employing abstract reasoning to come up with a natural language hypothesis, while being reminded that this should generalize to other, similar UUs. An example of abstraction being performed is shown in table 3.
Exploration Exploration as a strategy goes beyond attempting to reason about blind spots that result from discovered UUs. Instead, exploration focuses on using existing hypotheses to generalize to a new hypothesis and thus find additional, previously undiscovered blind spots to gain new knowledge regarding the model’s failure modes. Humans and LLMs are encouraged to use extrapolative thinking to form a hypothesis that they believe covers new potential failure modes. To this end, we request they ensure that the generated hypothesis is dissimilar compared to the already established hypothesis while targeting another potential weakness. For the exploration step, we present human-generated hypotheses to the human participants and vice versa. An example of exploration being performed is shown in table 4.
3.3 Synthetic Sample Generation
After generating hypotheses via abstraction or exploration, participants, whether human or LLM, are asked to generate synthetic samples, which we denote as $\bar{x}^{\prime}$, that follow these new hypotheses. These new samples must fit the overall structure of the used dataset, as well as be applicable to the overall context of the task (e.g., a movie-related sample for a movie review dataset). We thus obtain additional samples that should, according to our intuition, correspond to blind spots that the predictive model $\theta$ might have. Using these samples the training dataset is extended, resulting in $\mathcal{E}$, which consists of $\mathcal{D}$ and the additional generated samples $\bar{x}^{\prime}$ and their corresponding labels $\bar{y}^{\prime}$, which we also ask participants to provide. The generation procedure is the same, regardless of whether the used hypothesis was obtained via abstraction or exploration, and humans perform generation exclusively on human-generated hypotheses and vice versa.
Figure 2: Workflow: (a) Attacking the original model to obtain UUs; (b) Using these UUs to extend the training data via generalization (figure 1) and thus obtain a more robust model; (c) Evaluating this retrained model.
4 Experimental Setup
Initially, we provide an overview of the overall experimental setup, followed by a detailed description of the setup for our experiments, shown schematically in figure 2. Additionally to addressing natural blind spots (section 5), we performed a study on human-induced blind spots (appendix A), in which we show that one can create blind spots in a model that do not naturally occur when training on the data set, and subsequently successfully target its UUs via our approach, thus eliminating that blind spot. For further contextualization of the data we collect, we show a selection of data for the IMDB task. In table 3, we present an example of hypothesis generalization from a UU via abstraction and in table 4 an example of hypothesis generalization from a previous hypothesis by exploration.
Dataset Task #Classes #Train #Validation #Test
MRPC SE 2 3,668 408 1,725
IMDB SA 2 25,000 12,500 12,500
QNLI NLI 2 104,743 5,463 5,463
Table 2: Datasets used, including the task type, number of classes and number of samples in each of the test, validation, and training sets. Note the split of the original IMDB test set into validation and test set.
Original/Perturbed Samples O: […] Footage of busy highways, planes flying overhead, poor lightning effects, soldiers doing their thing, they’re all used in a [[haphazard]] manner, sometimes repeatedly, and they often only have a loose connection to the story Wood is trying to tell. […] P: […] Footage of busy highways, planes flying overhead, poor lightning effects, soldiers doing their thing, they’re all used in a [[thoughtless]] manner, sometimes repeatedly, and they often only have a loose connection to the story Wood is trying to tell. […]
Human Abstraction Hypothesis “Haphazard can be intentional, vs thoughtless is unplanned. It feels more of an unwillingness rather than a lack of ability”
LLM Abstraction Hypothesis “Hypothesis: The replacement of specific adjectives in the perturbed sample alters the sentiment and contributes to misclassification. Reasoning: The adjective "thoughtless" in the perturbed sample, replacing "haphazard" may change the perception of the director’s skills and use of stock footage. This alteration in sentiment could lead to a different classification outcome.”
Table 3: Examples of hypothesis generalization using abstraction for the IMDB dataset. Starting from a UU made up of an original and its corresponding perturbed sample, abstraction is performed by a human or LLM.
Human Abstraction Hypothesis “The pattern that may have caused the wrong prediction may be related to the substitution of the words with their synonym, for example ""bad"" with it’s synonym ""wicked"".”
Human Exploration Hypothesis “The hypothesis could be that the words were potentially swappped or replaced, such as ""bad"" been replaced by ""evil"" which could lead to an incorrect prediction.”
LLM Abstraction Hypothesis “The change in the word "appalling" to "horrifying" may have caused the misclassification. Both "appalling" and "horrifying" have negative connotations and can be used to describe something that is disturbing or shocking. However, the word "appalling" may be considered more extreme and may have a stronger negative connotation compared to "horrifying". This subtle difference in the intensity of the negative sentiment may have caused the misclassification by the algorithm.”
LLM Exploration Hypothesis “The use of the word "hypocrite" may have caused the misclassification. Reasoning: The word "hypocrite" implies that the person is saying one thing but doing another, which may be considered a negative trait. However, some people may not interpret the person in the sample as a hypocrite, leading to a difference in sentiment analysis. This difference in interpretation may have caused the misclassification by the algorithm.”
Table 4: Examples of hypothesis generalization using exploration for the IMDB dataset. Starting from a hypothesis obtained via abstraction, exploration is performed by a human or LLM.
4.1 Verification
For quality control purposes of the human contributions we include two attention checks in each survey to determine if crowdworkers are paying attention Oppenheimer et al. (2009). Additionally, we eliminate low-effort responses submitted by crowdworkers. Our goal for this is not to select high-quality responses, but instead to eliminate answers that were made in bad faith, e.g. by copy-pasting the same response repeatedly. For automated quality control for responses by humans and LLMs our approach is two-fold. For all generated text, whether hypothesis or sample, we ensure a certain character length ($\text{char}_{\text{min}}=40$) to exclude responses that have insufficient usable content. Finally, we employ an automatic quantitative evaluation technique in the form of BERTScore (Zhang et al., 2020). Here, we use the handcrafted hypothesis set and generations from appendix A as the reference set. If the score of a new hypothesis or sample fall below a threshold $S_{min}$, then we exclude it and repeat the previous step to obtain a replacement. In practice, we find that $S_{min}=0.5$ gives a good balance between accuracy and ensuring textual diversity.
4.2 Datasets and Models
We systematically evaluate the effectiveness of our approach on a range of distinct classification tasks, which were selected gauge the adaptability of our approach across different levels of task complexity. We focus on sentiment analysis (SA), semantic equivalence (SE), and natural language inference (NLI) for our experimental evaluation. To this end, we use the popular IMDB Maas et al. (2011) dataset for the SA task, the MRPC Dolan and Brockett (2005) dataset for the SE task, and the QNLI Rajpurkar et al. (2016) datasets for the NLI portion of our experiment. An overview of the details of the used datasets can be found in table 2. We choose this combination of datasets because they exhibit significant disparities in the training sample sizes, along with varying levels of task complexity. We finetune a BERT Devlin et al. (2019) model to perform the classification task and use the bert-base-uncased implementation provided by Wolf et al. (2020).
For blind spot mitigation, we restrict the number of hypotheses to 1% of the training set size for both abstraction-derived and exploration-derived hypotheses. Additionally, each generated sample corresponds to a single hypothesis, and we utilize each hypothesis, resulting in the number of new samples being equal to 2% of the training set size. These values can be effectively regarded as budget-dependent hyperparameters. The even split between abstraction-derived and exploration-derived hypotheses is also a choice that can be optimized, which we leave for future work. In line with previous works Lakkaraju et al. (2017); Bansal and Weld (2018); Han et al. (2021), we define $\tau=0.65$. For the generation of hypotheses or samples via LLM, we use GPT version gpt-3.5-turbo-1106 at a temperature $T=0.7$ with the default system prompt. All models are trained using a learning rate of $2\times 10^{-4}$ and a batch size of 64 for 10 epochs.
4.3 Perturbations
We assume no knowledge of the model parameters, i.e. a black-box setting, and focus on perturbation techniques that are most likely to maintain the same meaning post perturbation and thus preserve the underlying true label. We implement two state-of-the-art perturbation techniques via the TextAttack framework Morris et al. (2020), namely TextFooler Jin et al. (2020) and DeepWordBug Gao et al. (2018). TextFooler focuses on semantically and syntactically similar synonym swap perturbations. It identifies the most important words for the target model and then replaces them with semantically similar and grammatically correct words until the prediction is altered. DeepWordBug instead focuses on a range of character level perturbations that aim to minimize the edit distance of the perturbation. We thus cover two major lines of adversarial perturbations, i.e. world-level and character-level. Note that perturbation methods are not a necessity to uncover blind spots. Instead, we employ them to perform proactive instead of reactive illumination of blind spots and thus find additional instances failure modes.
4.4 Evaluation
We explore a range of metrics to evaluate the effectiveness of our approach. Specifically, we assess the accuracy of a model, the percentage of successful perturbations achieved by the perturbation technique, and the number of UUs that occur due to the perturbations. All of these are recorded for the model trained on the original dataset and for the one retrained on $\mathcal{E}$. The overall accuracy of the model on the test set gives basic insight into the general performance, while the successful perturbation percentage and the number of UUs indicate the overall robustness of the model and prevalence of blind spots that lead to high confidence misclassifications, respectively. It is desirable to maximize the first while minimizing the latter two. As an increasing number of blind spots are illuminated in the classification model, the number of UUs is reduced. As such, the model is less prone to catastrophic misclassifications due to perturbation attacks or other out-of-distribution data.
5 Natural Blind Spots
Unlike synthetic blind spots, which are induced via biasing, natural blind spots are present in the model due to natural biases in the train set. Beyond studying whether our approach is indeed successful in reducing blind spots across our diverse choice of classification tasks, we aim to compare human to LLM performance in terms of effectiveness, ability to scale, and ease of use.
One crucial difference in instruction between the human and LLM procedures is the number of examples given, with humans receiving two, fully worked out examples per participant, while no example is given to the LLM. This was done to lower the amount of guidance given to the model, as few-shot prompting the LLM leads to a highly uniform set of samples, even at higher temperature values. We thus intentionally introduce more variability, akin to what naturally transpires in the human procedure, where crowdworkers possess distinct perspectives and thoughts. The structure and content of the user study presented to the crowd workers is shown in appendix C, while the prompt structure used for the LLM is shown in appendix D.
TextFooler DeepWordBug Acc. (%) $\uparrow$ UUs (#) $\downarrow$ Acc. (%) $\uparrow$ UUs (#) $\downarrow$ MRPC Original 82.38 952 82.38 936 LLM Data 81.57 851 82.23 882 Human Data 81.58 418 82.10 802 IMDB Original 94.84 1882 95.40 1682 LLM Data 95.40 1241 94.41 1448 Human Data 94.43 1518 95.74 1412 QNLI Original 89.88 1923 89.88 2597 LLM Data 89.31 1536 89.38 1746 Human Data 89.42 2028 89.38 2325
Table 5: Results of natural blind spot study for accuracy, perturbation success rate, and number of UUs. Rows show results for the original and retrained models (using human- or LLM-generated synthetic samples) by dataset.
5.1 Natural Blind Spot Study Results
The results for the addition of the human- and LLM-generated samples are shown in table 5.
Accuracy and UUs Regarding accuracy, which serves as a proxy for general performance, we find that the procedure employed, whether human- or LLM-based, has no significant impact. Across the various datasets, the accuracy remains consistent, with a difference of no more than $\pm$ 1% between the original and retrained models. With regards to our main metric, the number of UUs, we observe that our method can successfully lead to decreases, with a maximum reduction of 56.09% for the human-based retraining on the MRPC dataset using TextFooler. The only exception to this is the human-based procedure on the QNLI dataset with the TextFooler attack, where we saw an increase of 5.46%. On average, across the two perturbation methods, there is a reduction in UUs of 19.54% and 16.80% for LLM-based and human-based retraining, respectively.
Perturbations A detailed overview of the perturbation statistics can be seen in appendix B. Perturbation success rate, which serves as a proxy for general model robustness against attacks, indicates that leveraging the LLM-based method leads to more stable results, with reductions of up to 8.82% for the IMDB dataset using DeepWordBug, while the human-based data leads to wider swings in performance.
Human Versus LLM Performance The decrease in UUs for LLMs is more substantial compared to the decrease observed as a result of the use of human-generated samples, as well as more predictable, suggesting that LLMs create quality hypotheses and samples more consistently. However, in the case of the TextFooler perturbation method on the MRPC dataset, the human-based approach exhibits a remarkable reduction in UUs, outperforming the LLM-based one significantly. This indicates a strong task dependency of the effectiveness of humans in characterizing blind spots using our proposed approach.
The task-dependent variability in human performance is evident in the undesired outcomes for the QNLI dataset when using TextFooler. Contrary to expectations, the responses generated by humans actually result in an increase in UUs. The observed variability in performance can likely be attributed to the intuitiveness of the SE and SA tasks, which contrasts with the complex NLI task. Determining the equivalence of two sentences or discerning the sentiment of a movie review aligns with the everyday experience of individuals. As a result, these tasks are perceived as less complex and more intuitive to crowdworkers. This underscores the task dependency of non-expert human abilities and highlights how human performance can vary based on the nature of the task at hand.
5.2 Distribution of Successful Perturbations
To visualize the effect of the retraining across the perturbed samples, we track the confidence at which predictions are made. The confidence plots across datasets are shown in figure 3 and figure 4 for the TextFooler and DeepWordBug perturbation methods, respectively. We can observe a reduction in high confidence predictions across the retrained models; specifically, predictions with confidence of above 90% are reduced as a result of retraining. Additionally, we observe similar results across perturbation methods, indicating that the performance of our approach is not dependent on the type of perturbation method employed.
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111100200200200200400400400400Confidence# of Instances
(a) MRPC Original
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111100200200200200400400400400Confidence
(b) MRPC Retrain LLM
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111100200200200200400400400400Confidence
(c) MRPC Retrain Human
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111002,00020002{,}0002 , $0004{,}$ $0006{,}$  [^0] Confidence# of Instances
(d) IMDB Original
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111002,00020002{,}0002 , $0004{,}$ $0006{,}$  [^0] Confidence
(e) IDMB Retrain LLM
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111002,00020002{,}0002 , $0004{,}$ $0006{,}$  [^0] Confidence
(f) IMDB Retrain Human
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111005005005005001,00010001{,}0001 , 0001,50015001{,}5001 , 500Confidence# of Instances
(g) QNLI Original
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111005005005005001,00010001{,}0001 , 0001,50015001{,}5001 , 500Confidence
(h) QNLI Retrain LLM
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111005005005005001,00010001{,}0001 , 0001,50015001{,}5001 , 500Confidence
(i) QNLI Retrain Human
Figure 3: Plots of successful perturbations for all datasets when using TextFooler, showing the distribution of the number of instances across confidence bins.
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111100100100100100200200200200300300300300Confidence# of Instances
(a) MRPC Original
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111100100100100100200200200200300300300300Confidence
(b) MRPC Retrain LLM
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9111100100100100100200200200200300300300300Confidence
(c) MRPC Retrain Human
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111002,00020002{,}0002 , $0004{,}$  [^0] Confidence# of Instances
(d) IMDB Original
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111002,00020002{,}0002 , $0004{,}$  [^0] Confidence
(e) IDMB Retrain LLM
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111002,00020002{,}0002 , $0004{,}$  [^0] Confidence
(f) IMDB Retrain Human
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111005005005005001,00010001{,}0001 , 0001,50015001{,}5001 , 500Confidence# of Instances
(g) QNLI Original
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111005005005005001,00010001{,}0001 , 0001,50015001{,}5001 , 500Confidence
(h) QNLI Retrain LLM
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111005005005005001,00010001{,}0001 , 0001,50015001{,}5001 , 500Confidence
(i) QNLI Retrain Human
Figure 4: Plots of successful perturbations for all datasets when using DeepWordBug, showing the distribution of the number of instances across confidence bins.
6 Discussion
Our results demonstrate that accuracy remains consistent regardless of whether the additional synthetic samples are generated by a human or LLM, while the perturbation success rate exhibits varying degrees of change. The LLM-based approach has a greater capacity for reducing UUs compared to human-generated samples. However, it is important to note that human-generated samples have the potential to be more effective in certain circumstances. Overall, our findings highlight the strengths and limitations of both LLM- and human-based approaches, and their potential synergies in improving model performance and robustness.
6.1 LLM Capabilities
In our investigation, we found that the quality of responses obtained from human participants surpasses those of LLM in a limited number of cases, particularly when higher-order thinking skills were necessary. For example, in one of our perturbed samples, we had the date “June 15” in a sample changed to “John 15”. A high-performing crowdworker was able to discern the relation to the Bible verse that was most likely the cause of the UU, while the LLM was not. While the number of such high-quality responses is small, when present, their impact on reducing UUs can be significant and compensates for the many low-quality responses. Thus, we posit that the quality ceiling for humans is higher, although LLMs more consistently provide acceptable hypotheses and samples. It is worth noting that the leading reasons for poor-quality responses are either insufficient understanding of the task or a lack of motivation among participants Gadiraju et al. (2017).
6.2 Scalability and Ease of Use
Turning to the practical aspects of our study, we obtain important insights regarding the costs and time required for conducting human and LLM generalization experiments. The human study, involving 168 participants, incurred a cost of \$1072 at an hourly rate of \$12, whereas the LLM experiment had a significantly lower cost of \$46 for the same number of generalizations and samples generated. Though difficult to quantify, the collection of the data from humans via surveys also takes significantly longer than collecting LLM data. This reinforces the fact that the LLM-based approach is significantly more scalable, as it is not only cheaper but also yields data almost instantaneously, while obtaining data from humans comes with a significant delay. Even though the LLM-based procedure is clearly the most scalable, there might be some high-stakes or specialist applications where a human-based or hybrid procedure is most useful. These findings shed light on the resource implications of conducting human and LLM characterization studies, thereby aiding researchers in planning and allocating resources effectively.
7 Conclusion
We propose an approach to identify and mitigate blind spots in NLP models through generalization by humans or LLMs, followed by synthetic sample generation to address UUs and thus enhance model robustness. Our evaluation shows that our approach is successful in achieving a significant reduction in UUs across datasets while maintaining accuracy. We find that LLMs outperform humans in characterizing blind spots and that there is a significant task dependency to the quality of human contributions.
8 Limitations
It should be noted that the chosen method of perturbation inevitably influences the hypotheses generated by human participants and LLMs alike. Since our chosen perturbation approaches aim to swap in synonyms for vulnerable words or modify characters, found blind spots will only be due to these particular perturbations, thus biasing the characterization and generation in favor of such errors. Another limitation is the potential fallibility of humans or LLMs. Leveraging humans for the generalization and generation tasks risks ignoring blind spots that are not identifiable by humans. The same can be said for LLMs, which might miss human-identifiable blind spots. Further, contributions made by humans may be biased themselves due to personal biases or motivations. We try to address this through the previously discussed verification step. Finally, one notable challenge we encountered was the difficulty of definitively proving the origin of the responses provided by a small minority of participants, as their responses appear to resemble those generated by ChatGPT when asked to perform the same task. This echos the findings by other researchers, that crowdworkers are increasingly using LLMs to generate their responses Veselovsky et al. (2023). While this is just an observation that we cannot definitively prove, it raises the concern that some participants may not have authored their responses themselves, undermining the integrity of our study.
Appendix A Synthetic Blind Spots
We use the synthetic blind spot study akin to a sanity check for our approach. As such, compared to the full natural blind spot study, we use a only a single task, a simpler model architecture, and make other simplifications to our mitigation process. We select an LSTM Hochreiter and Schmidhuber (1997) as our model of choice due to the absence of pretraining and apply the TextFooler perturbation method on the SA task. The LSTM used is the standard version of the Bi-LSTM provided by Morris et al. (2020).
Original Retrain
Accuracy (%) Perturbation (%) UUs (#) Accuracy (%) Perturbation (%) UUs (#)
Clean 88.03 82.22 1725 88.03 82.21 784
Biased R 78.55 78.56 3785 78.61 78.58 2593
Biased P 75.10 75.02 4607 74.25 73.12 1201
Biased N 76.64 76.64 4394 77.38 77.35 845
Biased PN 74.17 73.94 9231 74.81 74.01 2331
Table 6: Results of synthetic blind spot study for accuracy, perturbation success rate, and number of UUs before and after retraining for all LSTM model variants. The used perturbation method is TextFooler and the dataset is IMDB.
Blindspot Creation To assess whether our method can tackle existing synthetic blind spots we perform a type of Controlled Synthetic Data Check Nauta et al. (2023). We create synthetic blind spots by systematically excluding some data from training that have commonalities, namely containing a positive or negative term according to lexica by Liu et al. (2005). Here, we randomly subsample 600 of each as our selection of positive and negative terms, due to the extensive nature of the lexica.
We create a false positive blind spot by removing samples from the train set using our selection of negative terms, resulting in a negatively biased LSTM (N). Similarly, we create a false negative blind spot, resulting in a positively biased LSTM (P), as well as a blind spot resulting from a selection of 50% randomly chosen terms from each, leading to a positive/negative biased LSTM (PN). For comparison, we also include a randomly biased LSTM (R), where samples were removed from the train set randomly to obtain a size comparable to the P, N, and PN ones. [^2] Size of training sets: $N_{Clean}=25,000$, $N_{R}=2,500$, $N_{P}=2,439$, $N_{N}=3,138$, and $N_{PN}=2,438$.
Blind Spot Mitigation After creating the synthetic blind spots through biasing, the authors perform the generalization procedure and provide handcrafted hypotheses that precisely describe these, similar to golden labels. To generate the new samples from our handcrafted hypotheses, we prompt ChatGPT to generate movie review-related sentences (to fit the chosen task) that follow a given hypothesis. This was done in an attempt to simplify the procedure by taking advantage of human strengths, generalization and extrapolative thinking, and LLM strengths, low-cost text generation, simultaneously.
A.1 Synthetic Blind Spot Study Results
The mitigation results of this human-LLM approach for our Controlled Synthetic Data Check can be seen in table 6.
Outcomes of Biasing As can be seen in the first column of table 6, before retraining, the overall test accuracy declines in line with the degree to which the train set is biased. Interestingly, the percentage of successful perturbations by TextFooler, i.e., the percentage of successful label flips, closely follows the overall accuracy. This mirrors the findings of Tsipras et al. (2019), that there is a strong relationship between high accuracy and brittleness – or a lack of robustness. The number of occurring UUs as a result of the perturbation does not follow this trend, instead increasing as the training data becomes more biased, as expected. This poses an interesting optimization problem since the model becomes most robust in general terms, i.e., the successful perturbation percentage falls, but simultaneously there is a significant uptick in blind spots as the training sets become more biased.
Blind Spot Mitigation The effect of retraining on the overall accuracy and perturbation success rate is minimal, with accuracy changing by no more than $\pm$ 1% and perturbation success rate changing no more than $\pm$ 2%. However, the number of found UUs decreases drastically due to the retraining, with reductions of 73.93%, 80.77%, and 74.75% for the biased P, N, and PN models, respectively. The clean and randomly biased models also show a reduction, though less significant at 54.55% and 31.49%, respectively. These results confirm that our method can be used to target synthetic blind spots found in biased models through the use of hypotheses and generated instances, without significantly affecting the performance or general robustness of the model.
Appendix B Perturbation Statistics
To add additional context to the perturbation performed, we supply the detailed attack statistics across all performed perturbations. Specifically, we report Original Accuracy and Accuracy Under Attack are reported, which are the classifier accuracy on its own and while under attack. Further, Attack Success Rate is shown, which is the percentage of successful perturbation attempts to failed ones. Finally, we report the number of Perturbed Words, the percentage of words that are perturbed, the Words per Input, the average number of words per input, and the Average Number of Queries, which is how many tries it took the perturbation method to find the best attack. The attack statistics for TextFooler attacks are shown in table 7 while the ones for DeepWordBug attacks are shown in table 8.
[…]
Original Accuracy (%) 82.38 81.57 81.58 94.84 95.40 94.43 89.88 89.31 89.42
Accuracy Under Attack (%) 9.80 17.40 12.99 10.18 10.44 19.21 8.91 11.67 14.89
Attack Success Rate (%) 71.83 64.87 68.29 88.46 93.18 63.85 87.35 86.80 78.84
Perturbed Words (%) 7.70 9.9 8.51 4.59 7.62 9.02 6.12 8.80 9.57
Words per Input 39.3 39.3 39.3 230.0 230.0 230.0 37.9 37.9 37.9
Avg. Number of Queries 51.40 68.62 55.17 185.24 184.94 198.31 49.38 51.27 56.11
Table 7: Perturbation statistics across datasets and models for attacks with TextFooler. Subscripts O, L, and H denote the original, LLM-retrained, and human-retrained models, respectively.
[…]
Original Accuracy (%) 82.38 82.23 82.10 95.40 95.41 95.74 89.88 89.38 89.38
Accuracy Under Attack (%) 7.78 13.73 11.94 9.54 21.43 15.32 8.21 9.90 7.30
Attack Success Rate (%) 72.00 70.38 72.64 59.41 50.59 79.70 77.54 79.74 82.08
Perturbed Words (%) 8.47 9.18 9.03 6.43 8.11 13.09 7.99 8.32 11.03
Words per Input 39.3 39.3 39.3 230.0 230.0 230.0 37.9 37.9 37.9
Avg. Number of Queries 56.92 64.37 58.61 199.32 211.65 201.44 34.91 33.53 49.09
Table 8: Perturbation statistics across datasets and models for attacks with DeepWordBug. Subscripts O, L, and H denote the original, LLM-retrained, and human-retrained models, respectively.
Appendix C User Study Structure
We use Prolific as a crowdsourcing platform for all our participants. Below, we present the structure followed by all survey participants for the generalization user study, consisting of an initial disclaimer, an instruction set, examples, and finally the questions. Here, we use the abstraction and exploration assignments on the IMDB dataset as an example. The workflow is very similar between the different generalization assignments and datasets (MRPC, IMDB, or QNLI), with only slight differences in the wording between the surveys to fit the task and dataset used, as they all present the crowd worker with some input and result in plain text output. For the generation assignment, crowdworkers are asked to perform the same steps, with relevant examples related to the structure of the dataset being shown, before finally contributing usable samples based on shown hypotheses.
C.1 Abstraction on IMDB
Disclaimer Crowdworkers were shown an initial disclaimer to inform them that our governing ethics body sanctions this survey and to remind them not to share personal information:
  • “Welcome to the Hypothesis Exploration Survey! Please carefully read the following: You are invited to participate in our research study. This study is fully sanctioned by our governing ethics body, as is the handling and storing of the resulting data. This research study aims to use your creativity and generalization ability to come up with new abstractions. It will take you approximately 25 minutes to complete. As with any online activity, the risk of a breach is always possible. To the best of our ability, your answers in this study will remain confidential. We will minimize any risks by making this survey completely anonymous. Therefore, please do not provide any personal information anywhere. The anonymous results might be shared publicly in the future. Participation in this study is entirely voluntary, and you can withdraw anytime. Feel free to contact us with any questions or feedback you might have.”
Instructions Crowdworkers were then introduced to the specific task (SE, SA, or NLI) as follows:
  • “Please read the following examples carefully. All tasks in this survey are related to a single task, sentiment analysis, which tests the sentiment of a sentence is either positive or negative, applied to movie reviews. The goal here is to use your creativity and ability to generalize to spot patterns and come up with new possible samples. A fully worked-out example can be found below, with user-generated text, similar to what you are expected to write, in italic and instructions bold. You will receive all relevant instructions again when for each question.”
Examples Then, they were presented with two examples that match the dataset used, as well as the task (abstraction, expansion, or generation), before being asked if they understood the examples:
  • There is a sentence pair below, with one original sample (O) and a perturbed one (P), which is similar but had some things changed (shown in double square brackets). These changes may relate to a pattern, related to semantics, syntax, specific words, or something else in the samples, that leads to the wrong True or False label being predicted for semantic similarity.
  • Example 1 – The two samples are:
    O: There was an overarching [[story]] that was [[refusing]] to reveal itself to me. P: There was an overarching [[narrative]] that was [[unable]] to reveal itself to me.
    Formulate a hypothesis on what this pattern for O and P might be and enter it below. Try to be specific when formulating a hypothesis.
    The pattern that caused the wrong prediction may be related to the substitution of the word ""story"" with its synonym ""narrative"".
  • Example 2 – The two samples are:
    O: Overall, I [[loved]] the cinematography of this through and [[through]]. P: Overall, I [[looved]] the cinematography of this through and [[thr0ugh]].
    Formulate a hypothesis on what this pattern for O and P might be and enter it below. Try to be specific when formulating a hypothesis.
    Several words have been misspelled in the samples, all related to the letter ""o"". Either more letters are added ""oo"" or the letter is substituted with a number ""0"" that looks similar, making it easy to misread.
Main Questions Finally, the actual questions preceding the text entry field used for data collection all have the same structure with the unique O and P sentences substituted in for each question:
  • “The two samples are:
    O: <original sentence> P: <perturbed sentence>
    Formulate a hypothesis on what this pattern might be and enter it below. Try to be specific when formulating a hypothesis.
Appendix D Used LLM Prompts
We specifically instruct the LLM to split its hypothesis from its reasoning because, in our experience, this leads to a clearer and more useful answer for further steps.
D.1 Abstraction Prompt
  • “There is a sentence pair below, with one original sample (O) and a perturbed one (P), which is similar but had some things changed. These changes may relate to a pattern, related to semantics, syntax, specific words, or something else in the samples, that leads to them being the reason the sample is misclassified by a classification algorithm. This misclassification is made at a high level of confidence.
    The model is not trained on the two samples. The two samples relate to {task} and are:
    O: {sentence[0]}
    P: {sentence[1]}
    Formulate a hypothesis on what this pattern might be. Try to be specific when formulating a hypothesis. Your response should always follow the format:
    Hypothesis: <<hypothesis>>
    Reasoning: <<reasoning>>”
D.2 Exploration Prompt
  • “There is a sentence pair, with one original sample (O) and a perturbed one (P), which is similar but had some things changed. These changes may relate to a pattern, related to semantics, syntax, specific words, or something else, that leads to them being the reason the sample is misclassified by a classification algorithm. This misclassification is made at a high level of confidence.
    The model is not trained on the two samples. The two samples relate to {task}
    There is an existing hypothesis regarding the samples, that may capture a pattern related to semantics, syntax, specific words, or something else in the sample pair. This pattern leads to a misclassification of the sample.
    The hypothesis is: {hypothesis}
    Formulate a new hypothesis regarding those sentence samples that is concerned with the same topic but is applied to a different possible pattern that could also lead to a misclassification. Try to be specific when formulating a new hypothesis. Your response should always follow the format:
    Hypothesis: <<hypothesis>>
    Reasoning: <<reasoning>>”
D.3 Generation Prompt
  • “There is a sentence pair, with one original sample (O) and a perturbed one (P), which is similar but had some things changed. These changes may relate to a pattern, related to semantics, syntax, specific words, or something else, that leads to them being the reason the sample is misclassified by a classification algorithm. This misclassification is made at a high level of confidence.
    The model is not trained on the two samples.
    A hypothesis has been formulated regarding the samples, that may capture a pattern related to semantics, syntax, specific words, or something else in the sample pair. These samples led to a classification algorithm misclassifying them at a high level of confidence.
    Given the samples and a previously generalized hypothesis, generate one new sample made up of one or more sentences that relate to {task} and could have a similar effect on the classification algorithm.
    The new sample should be varied and detailed. Follow the logic laid out in the given hypothesis and follow the format of the sample pair (O and P) exactly. Also include whether the new sample should be given a <<positive>> or <<negative>> label for the task: {task}.
    The hypothesis is: {hypothesis}
    Your response should always follow the format:
    Sample: <<sample>>
    Label: <<label>>
    Reasoning: <<reasoning>>”